OpenStudy (anonymous):

9(x + 2) = 3(x - 2) x = 3(x - 6) + 2 = 4(x + 2) - 21 x =

OpenStudy (anonymous):

First distribute the outside number. 9(x + 2) = 3(x - 2) 9x + 18 = 3x - 6 Then get the variable by itself. 9x + 18 = 3x - 6 6x + 18 = -6 6x = -24 x= -4 Now to plug it in: 9(-4 + 2) = 3(-4 - 2) 9(-2) = 3(-6) -18 = -18 Therefore, for problem 1, x = -4